OpenStudy (123panda):

This is NOT a multiple choice question, and I need help super bad! Please help!!

OpenStudy (chrishogan):

Sure, what's up?

OpenStudy (mathmale):

Save time by posting your question up front, immediately.

OpenStudy (123panda):

This is the question

OpenStudy (123panda):

There is also this question

OpenStudy (mathmale):

One question at a time, please. First situation: f(x)=x^2+2 and g(x)=2x^2+5 You want to find the composition g( f(x) ). This means: let the function f(x) become the INPUT to the function g(x). Start with g(x): g(x) = (x)^2 + 2. Replace the first x with 'f(x)' and the second x with 2x^2+5. Show your work, please.